OpenStudy (anonymous):

solve for x, b(5px-3c)+a(qx-4)

OpenStudy (anonymous):

Distribute. Where is your equal sign?

OpenStudy (luigi0210):

^I was gonna say the same thing

OpenStudy (luigi0210):

This is an expression not an equation so we can't solve, just simplify

OpenStudy (anonymous):

@Luigi0210 @gypsy1274 b(5px-3c)=a(qx-4)

OpenStudy (luigi0210):

Then yea distirbute first

OpenStudy (luigi0210):

*distribute

OpenStudy (anonymous):

Much better. That shift button can be awfully shifty...

OpenStudy (anonymous):

ok so then you get b5px-3bc+aqx-4a

OpenStudy (luigi0210):

@gypsy1274 's got you

OpenStudy (anonymous):

There you go with that shifty shift key again.... 5bpx-3bc=aqx-4a Now use addition and subtraction to get the x terms on the left and the non x terms on the right.

OpenStudy (anonymous):

@Luigi0210 don't go too far...Your advice is always appreciated.

OpenStudy (anonymous):

okk so you get \[\frac{ aqx-4a+3bc }{ 5bp}\] ?

OpenStudy (anonymous):

wait thats not right because theres another x hmm

OpenStudy (anonymous):

One step at a time. 5bpx-3bc=aqx-4a Pick one term to work with first. Then decide what you need to do to get it where you want it.

OpenStudy (anonymous):

so what do i do first? im honestly so confused

OpenStudy (anonymous):

You started here: \(b(5px-3c)=a(qx-4)\) Then distributed to get: \(5bpx-3bc=aqx-4a\) The next step is to get your x terms to the left side.

OpenStudy (anonymous):

so divide by aqx?

OpenStudy (anonymous):

\(aqx\) is not multiplied by anything, so don't divide it. I would subtract it from both sides.

OpenStudy (anonymous):

so the answer is \[\frac{ 4a+3bc }{ 5bp-aq }\]

OpenStudy (anonymous):

One of us dropped a negative sign on the floor, re-checking now...

OpenStudy (anonymous):

I'm pretty sure it was you....look around the desk and floor, I'm sure you will find it. :-) the \(4a\) was negative and should remain negative.

OpenStudy (anonymous):

op i found it! it rolled under my bed not sure how it got there hmm haha thanks :)

OpenStudy (anonymous):

I have that problem too. Those negatives are very slippery little suckers....